OpenStudy (anonymous):

I need help.. ;-; A line has a slope of 9/4 and a y-intercept of −14. Which equations represent the line?

OpenStudy (anonymous):

@iGreen?

Nnesha (nnesha):

ok so slope and y-intercept form equation look like this y =mx +b where m is slope and b y-intercept so just plug that value where m and b are in the equation

Nnesha (nnesha):

do you understand ??

OpenStudy (anonymous):

I think so, I got y=9/4x+-14. But you have to choose all the answers that are correct...

Nnesha (nnesha):

ok so positive and negative = what

Nnesha (nnesha):

positive and negative become negative so y = 9/4x -14